In an alloy of 150 pounds of zinc and copper, there are 100 pounds of copper. How much copper must be added so that the alloy may be 10% zinc?
Found 2 solutions by Boreal, stanbon:
Answer by Boreal(15235)

The fraction is Zn/Cu+Zn
right now it is 50/150, if 150 pounds of both and 100 pounds of Cu, the rest has to be Zn. Let x=pounds of copper added
50/(100+x+50)=0.1
50/(150+x)=0.1
50=15+0.1x
35=0.1x
x=350 pounds of copper
At that point, the alloy will weigh 500 pounds, 50 of which will be zinc.

Answer by stanbon(75887)

In an alloy of 150 pounds of zinc and copper, there are 100 pounds of copper. How much copper must be added so that the alloy may be 10% zinc?
------
Equation::
zinc + zinc = zinc
----
(1/3)150 + 0*x = 0.10(150+x)
-------
50 + 0 = 15 + 0.1x
35 = 0.1x
x = 350 lb (amt of copper to add)
---------------
Cheers,
Stan H.
------------